day 10-June-99, update

Marj Corcoran (corcoran@physics.rice.edu)
Thu, 10 Jun 1999 15:50:49 -0500 (CDT)

Shift Summary for DAY shift, 10-June-99

On shift: S. Glazov, M. Corcoran

Also present: A cast of thousands....Ed, Bud, Rick K, Peter Shanahan,
Jim Graham, Brian LaVoy, Elizabeth, Ashkan, Sydney, Julie, Jason L.
Doug, Viv, Dave N.

Hours of beam delivered at all intensitites: 0
Hours of beam delivered at nominal intensity:0
Hours of E832 data (special runs not included):0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: No beam all shift

==================================================================

Runs ended this shift:None

Tapes brought to FCC: None

===================================================================

Description of shift:

Many activities in the hall, beam is off for accelerator repairs.

Surveyors arrive at 8AM to realign DC2. Beam is turned off, interlocks
dropped. Surveyors are done by 11AM, they are able to realign DC2 as
well as it was originally, but it is still 0.5mm off at one corner.
Doug finds that the z position of DC2 has changed by 1.1mm, from
44.6622m to 44.6633m. DC1 and the regenerator z-positions are unchanged.

Bud moves in UPS for the vacuum system, Doug activates it.

Discriminator which feeds repeater near laser rack was replaced.
The output pulse from the old discriminator was too short, speculation is that
this was the problem causing hang-ups of this repeater.

Ed and Julie replace divider for CsI channel 2877. A more serious problem
is found with channel 3031. This channel had low occupancy for electron
calibration. Turns out the flange had fallen off the crystal! It was
reglued around 12:40, need to wait a few hours before reinstalling the tube.

Rick C. put the accidental counter back to the 1/2" hole.

Elizabeth adjusts VV' HV's and delays; also adds another trigger source
at lower threshold for the MA and moved CA3 channel to a different
shaper/splitter.

Laser light level increased by Peter Shanahan.

Ashkan removes bypass modules from CAMAC crates in stations 1,2,3.

NM3S power supply repaired by ops.

Discovered that the DVM used to check chamber voltages during plateaus
last night had a low battery and may have been reading the voltage wrong by
a few percent. The plateaus done last night are suspect. Replacing the battery
fixed the problem with the DVM.

Alcohol added to the bubbler by DN, level now 3.5".

Toward end of shift, Elizabeth notices that the regenerator is making funny
sounds...Doug is paged.

===================================================================

Followup notes on previous shift

===================================================================
Items to follow up on next shift

CsI channel 3031 glue still drying. When glue is dry enough, the
phototube needs to be reseated.

Regenerator mover may have a problem, needs to be checked.

Probably want to redo the chamber plateaus due to problem with DVM.